# DRT: Deep Reasoning Translation via Long Chain-of-Thought

### v3

@@ Abstract
Recently, O1-like models have emerged as representative examples, illustrating the effectiveness of long chain-of-thought (CoT) in reasoning tasks such as math and coding tasks. In this paper, we introduce DRT, an attempt to bring the success of long CoT to neural machine translation (MT). Specifically, in view of the literature books that might involve similes and metaphors, translating these texts to a target language is very difficult in practice due to cultural differences. In such cases, literal translation often fails to convey the intended meaning effectively. Even for professional human translators, considerable thought must be given to preserving semantics throughout the translation process. To simulate LLMs’ long thought ability in MT, we first mine sentences containing similes or metaphors from existing literature books, and then develop a multi-agent framework to translate these sentences via long thought. In the multi-agent framework, a translator is used to iteratively translate the source sentence under the suggestions provided by an advisor. To ensure the effectiveness of the long thoughts, an evaluator is also employed to quantify the translation quality in each round. In this way, we collect tens of thousands of long-thought MT data, which is used to train our DRT. Using Qwen2.5 and LLama-3.1 as the backbones, DRT models can learn the thought process during machine translation, and outperform vanilla LLMs as well as LLMs which are simply fine-tuning on the paired sentences without long thought, showing its effectiveness.
@@
Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com [Uncaptioned image]DRT[Uncaptioned image]DRT-7B[Uncaptioned image]DRT-8B[Uncaptioned image]DRT-14B
@@ 1 Introduction
i) A suitable translation scenario to employ long thought in MT: Not all scenarios require long chain-of-thought (CoT)111“long CoT” is equal to “long thought”, and we alternatively use these two terms in this paper. during translation. For example, in simple expressions, literal translation can meet most needs, and translation via long CoT may be unnecessary. Inappropriate scenarios might cause the overthinking issue Chen et al. (2024).
@@ 2.3 Long Thought Reformulation
where $\mathcal{P}(s)$ denotes the multi-agent thought process for $s$, and $m$ is the number of iteration steps. To emphasize the valid thought process, translations without score change will be removed. That is, if $s^{i}$ is equal to $s^{i-1}$ ($i=1,2,...,m$), we will discard $\langle t^{i},f^{i},s^{i}\rangle$ in $\mathcal{P}(s)$, resulting in:
@@ 2.5 Quality Analyses
DRT-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
DRT-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
DRT-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
Table 3: Experimental results on literature translation. The bold and the underline denote the best and second-best performances, respectively. $\dagger$” and“$\ddagger$” denote statistically significant better than the corresponding SFT LLMs (w/o CoT) with t-test p < 0.01 and 0.05, respectively.
@@ 3.5 Inference Time Analysis
And if their words had taken another turn…if he himself had been less fastidious about intruding on another man’s secrets…it was cruel to think how thin a film had shut out rescue from all this guilt and misery. {CJK}UTF8gbsn 而如果他们的言辞转向了另一个方向……如果他自己对侵扰他人秘密不那么介意……想到正是那层薄薄的隔阂将救援拒之门外,使得这一切的罪恶和痛苦无法避免,真是残酷。 {CJK}UTF8gbsn 而且,假如他们的谈话走向不同……假如他自己不对干涉他人私事如此谨慎……那时他会意识到,只要一层薄薄的屏障就阻隔了救援,使得所有这些罪恶和痛苦无法解脱,这真是残忍。 {CJK}UTF8gbsn 如果他们的言辞稍有不同,他不那么介意侵扰别人的隐私,想到那层薄纱如何隔绝了救援,真是残忍。 {CJK}UTF8gbsn 若他们的言辞稍作转变……若他本人不那么介意侵入他人的心扉……想到这层薄薄的隔阂竟将所有内疚与痛苦隔离于救援之外,实为残忍。
This cold officer upon a monument, who dropped epithets unconcernedly down, would be finer as a dead man, he thought. {CJK}UTF8gbsn 他认为,这个站在纪念碑上的冷漠官员,若死了会更好他不带任何感情地抛下了一些称呼 {CJK}UTF8gbsn 这个冷冰冰的官员站在纪念碑上,毫不在意地抛下一些称号,他想,如果作为一个死人会更出色 {CJK}UTF8gbsn 这位冷酷的官员站在纪念碑上,随意地抛出贬义词,他觉得,作为死人会更显尊贵 {CJK}UTF8gbsn 纪念碑上的冷淡官员,漫不经心地吟咏那些修饰语,他心想,若化为亡者,或许更显尊贵。
I see on your cheek two tears which I know are hot as two sparks, and salt as two crystals of the sea. {CJK}UTF8gbsn 我看见你脸颊上有两滴泪水,我知道它们像两簇火花一样滚烫,像两颗海水结晶一样咸。 {CJK}UTF8gbsn 我见你面颊两颗泪,温热如火花咸涩如海晶 {CJK}UTF8gbsn 我见你脸颊上两颗泪珠,热如火花,咸似海盐。 {CJK}UTF8gbsn 我看见你脸颊上的两滴眼泪,它们既如两簇火花般炽热又似两粒海盐般咸涩
The night wore out, and, as he stood upon the bridge listening to the water as it splashed the river-walls of the Island of Paris, where the picturesque confusion of houses and cathedral shone bright in the light of the moon, the day came coldly, looking like a dead face out of the sky. {CJK}UTF8gbsn 夜深了,当他站在桥上,听着水花拍打巴黎岛河岸的声音,月光下房屋与教堂错落有致的景象闪耀着光芒时,冷冰冰的白天来临了,仿佛是从天空中露出的一张死人脸 {CJK}UTF8gbsn 夜晚消逝了,当他站在桥上,聆听水流拍打着巴黎岛的河岸,那里的房屋与大教堂构成了一幅混乱而又美丽的图景,在月光下闪耀着明亮的光芒,而白天则冷酷地到来,仿佛是天空中一张死人脸 {CJK}UTF8gbsn 夜幕渐渐消散,他站在桥上,聆听水花轻拍巴黎岛的河岸。月光下,岛上房屋与大教堂交织成一幅迷人的画卷,显得格外明亮。清晨悄然而至宛如从天际浮现的一张冷峻的脸庞。 {CJK}UTF8gbsn 夜色渐渐消逝,他站在桥上,倾听河水轻拍巴黎城岛的河岸。月光下,房屋与大教堂错落交织闪耀着迷人的光辉。白日冷冰冰地降临,宛如从天际浮现的一张死寂的
@@ A.1 Prompts in Literature Book Mining
SYSTEM PROMPT: You are assigned to translate an English literary work into Chinese. The text may include descriptions or expressions that embody English cultural nuances, which may not resonate with Chinese language habits. In such instances, a literal translation may not be appropriate; instead, these sentences should be paraphrased to convey their intended meaning effectively. USER PROMPT: The English sentence is provided as follows: <english sentence> {sentence} </english sentence> Please begin by assessing whether the English sentence contains any metaphors or similes. If there are none, respond with "no metaphors and no similes." If the English sentence does contain metaphors or similes, provide a literal translation of them, and then evaluate whether the literal translation is appropriate and easy for Chinese natives to understand. If it is suitable, format your response as follows (two lines): "your literal translation for metaphors/similes here (in Chinese)" "suitable" If it is unsuitable, please provide the reason for the unsuitability. Format your response as follows (three lines): "your literal translation for metaphors/similes here (in Chinese)" "unsuitable" "reason for unsuitability here (in Chinese)"
@@ A.2 Prompts in Multi-Agent Framework
Given an English sentence, identify the important words (usually nouns, verbs, technical terms, and named entities that require special attention in translation) and translate them into Chinese. Output the translations in JSON format, for example: {"EnglishWord1": "ChineseTranslation", "EnglishWord2": "ChineseTranslation"} The Chinese translations can be a single translation or multiple options as deemed appropriate.
@@
SYSTEM PROMPT: Given an English sentence and a JSON object containing potential translations of important keywords, produce a Chinese literal translation of the entire sentence. Please directly output the Chinese translation without any descriptions. USER PROMPT: <English Sentence> {sentence} </English Sentence> <Potential Keyword Translation> {keyword translation} </Potential Keyword Translation>
@@
SYSTEM PROMPT: Please evaluate the following Chinese translation of an English text. Rate the translation on a scale of 0 to 100, where: - 10 points: Poor translation; the text is somewhat understandable but contains significant errors and awkward phrasing that greatly hinder comprehension for a Chinese reader. - 30 points: Fair translation; the text conveys the basic meaning but lacks fluency and contains several awkward phrases or inaccuracies, making it challenging for a Chinese reader to fully grasp the intended message. - 50 points: Good translation; the text is mostly fluent and conveys the original meaning well, but may have minor awkwardness or slight inaccuracies that could confuse a Chinese reader. - 70 points: Very good translation; the text is smooth and natural, effectively conveying the intended meaning, but may still have minor issues that could slightly affect understanding for a Chinese reader. - 90 points: Excellent translation; the text is fluent and natural, conveying the original meaning clearly and effectively, with no significant issues that would hinder understanding for a Chinese reader. Please provide the reason first, followed by a score. Format your evaluation in the JSON structure below: {"reason": "reason for the score", "score": int}
@@ A.3 Prompts in Thought Reformulation
A student is engaged in the task of translating an English sentence into Chinese. The English sentence is as follows: <English Sentence> {sentence} </English Sentence> This student constantly thinks about and optimizes his translation. The whole process is shown as follows: <Translation Process> {translation process} </Translation Process> Please polish the whole translation process into a long first-person self-reflection description (use the present tense). The self-reflection should begin with selecting the keywords from the English sentence, translating the keywords, and then attempt to translate the whole sentence, and then think about whether the translation is good or not, and iteratively make translation attempts. Finally, make a final translation decision. Output the self-reflection description directly without any additional descriptions or explanations. Each line in the self-reflection description can be regarded as a reasoning step to the translation.
@@ Appendix C GPT-4o Evaluator
Score the following translation from English to Chinese with respect to the human reference on a continuous scale from 0 to 100, where score of zero means "no meaning preserved" and score of one hundred means "perfect preservation of meaning, with faithfulness, expressiveness, and elegance". English source: {src} Chinese human reference: {ref} Chinese translation: {hyp} Score:
@@
Score the following translation from English to Chinese on a continuous scale from 0 to 100, where score of zero means "no meaning preserved" and score of one hundred means "perfect preservation of meaning, with faithfulness, expressiveness, and elegance". English source: {src} Chinese translation: {hyp} Score:
@@
DRT-8B (Backbone: ${}_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
DRT-7B (Backbone: ${}_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
DRT-14B (Backbone: ${}_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
@@
Training Details. Llama-Factory Zheng et al. (2024) is used to instruct-tune LLMs. All LLMs are tuned on 8$\times$×NVIDIA A100 GPUs (40G) with 1e-5 learning rate and 8 (8$\times$1) batch size. We use the DeepSpeed ZeRO-3 optimization Rasley et al. (2020). Following Qin et al. (2024), we set the number of training epochs to 3, and the training process costs 70 GPU hours and 124 GPU hours for 7B and 14B models, respectively.

### v4

@@ Abstract
Recently, O1-like models have emerged as representative examples, illustrating the effectiveness of long chain-of-thought (CoT) in reasoning tasks such as math and coding tasks. In this paper, we introduce DRT, an attempt to bring the success of long CoT to neural machine translation (MT). Specifically, in view of the literature books that might involve similes and metaphors, translating these texts to a target language is very difficult in practice due to cultural differences. In such cases, literal translation often fails to convey the intended meaning effectively. Even for professional human translators, considerable thought must be given to preserving semantics throughout the translation process. To simulate LLMs’ long thought ability in MT, we first mine sentences containing similes or metaphors from existing literature books, and then develop a multi-agent framework to translate these sentences via long thought. In the multi-agent framework, a translator is used to iteratively translate the source sentence under the suggestions provided by an advisor. To ensure the effectiveness of the long thoughts, an evaluator is also employed to quantify the translation quality in each round. In this way, we collect tens of thousands of long-thought MT data, which is used to train our DRT. Using Qwen2.5 and LLama-3.1 as the backbones, DRT models can learn the thought process during machine translation, and outperform vanilla LLMs as well as LLMs which are simply fine-tuning on the paired sentences without long thought, showing its effectiveness. [^1] The synthesized data and model checkpoints are released at https://github.com/krystalan/DRT.
@@
Jiaan Wang,  Fandong Mengthanks:   Corresponding author.,  Yunlong Liang,  Jie Zhou Pattern Recognition Center, WeChat AI, Tencent Inc {torchwang,fandongmeng,yunlonliang,withtomzhou}@tencent.com
@@ 1 Introduction
i) A suitable translation scenario to employ long thought in MT: Not all scenarios require long chain-of-thought (CoT)222“long CoT” is equal to “long thought”, and we alternatively use these two terms in this paper. during translation. For example, in simple expressions, literal translation can meet most needs, and translation via long CoT may be unnecessary. Inappropriate scenarios might cause the overthinking issue Chen et al. (2024).
@@ 2.3 Long Thought Reformulation
where $\mathcal{P}(s)$ denotes the multi-agent thought process for $s$, and $m$ is the number of iteration steps. To emphasize the valid thought process, translations without score change will be removed. That is, if $s^{i}$ is equal to $s^{i-1}$ (i=1,2,,mi=1,2,...,m), we will discard $\langle t^{i},f^{i},s^{i}\rangle$ in $\mathcal{P}(s)$, resulting in:
@@ 2.5 Quality Analyses
DRT-8B $_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
DRT-7B $_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
DRT-14B $_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
[…]
@@ 3.5 Inference Time Analysis
And if their words had taken another turn…if he himself had been less fastidious about intruding on another man’s secrets…it was cruel to think how thin a film had shut out rescue from all this guilt and misery. 而如果他们的言辞转向了另一个方向……如果他自己对侵扰他人秘密不那么介意……想到正是那层薄薄的隔阂将救援拒之门外,使得这一切的罪恶和痛苦无法避免,真是残酷。 而且,假如他们的谈话走向不同……假如他自己不对干涉他人私事如此谨慎……那时他会意识到,只要一层薄薄的屏障就阻隔了救援,使得所有这些罪恶和痛苦无法解脱,这真是残忍。 如果他们的言辞稍有不同,他不那么介意侵扰别人的隐私,想到那层薄纱如何隔绝了救援,真是残忍。 若他们的言辞稍作转变……若他本人不那么介意侵入他人的心扉……想到这层薄薄的隔阂竟将所有内疚与痛苦隔离于救援之外,实为残忍。
This cold officer upon a monument, who dropped epithets unconcernedly down, would be finer as a dead man, he thought. 他认为,这个站在纪念碑上的冷漠官员,若死了会更好他不带任何感情地抛下了一些称呼 这个冷冰冰的官员站在纪念碑上,毫不在意地抛下一些称号,他想,如果作为一个死人会更出色 这位冷酷的官员站在纪念碑上,随意地抛出贬义词,他觉得,作为死人会更显尊贵 纪念碑上的冷淡官员,漫不经心地吟咏那些修饰语,他心想,若化为亡者,或许更显尊贵。
I see on your cheek two tears which I know are hot as two sparks, and salt as two crystals of the sea. 我看见你脸颊上有两滴泪水,我知道它们像两簇火花一样滚烫,像两颗海水结晶一样咸。 我见你面颊两颗泪,温热如火花咸涩如海晶 我见你脸颊上两颗泪珠,热如火花,咸似海盐。 我看见你脸颊上的两滴眼泪,它们既如两簇火花般炽热又似两粒海盐般咸涩
The night wore out, and, as he stood upon the bridge listening to the water as it splashed the river-walls of the Island of Paris, where the picturesque confusion of houses and cathedral shone bright in the light of the moon, the day came coldly, looking like a dead face out of the sky. 夜深了,当他站在桥上,听着水花拍打巴黎岛河岸的声音,月光下房屋与教堂错落有致的景象闪耀着光芒时,冷冰冰的白天来临了,仿佛是从天空中露出的一张死人脸 夜晚消逝了,当他站在桥上,聆听水流拍打着巴黎岛的河岸,那里的房屋与大教堂构成了一幅混乱而又美丽的图景,在月光下闪耀着明亮的光芒,而白天则冷酷地到来,仿佛是天空中一张死人脸 夜幕渐渐消散,他站在桥上,聆听水花轻拍巴黎岛的河岸。月光下,岛上房屋与大教堂交织成一幅迷人的画卷,显得格外明亮。清晨悄然而至宛如从天际浮现的一张冷峻的脸庞。 夜色渐渐消逝,他站在桥上,倾听河水轻拍巴黎城岛的河岸。月光下,房屋与大教堂错落交织闪耀着迷人的光辉。白日冷冰冰地降临,宛如从天际浮现的一张死寂的
@@ A.1 Prompts in Literature Book Mining
SYSTEM PROMPT: You are assigned to translate an English literary work into Chinese. The text may include descriptions or expressions that embody English cultural nuances, which may not resonate with Chinese language habits. In such instances, a literal translation may not be appropriate; instead, these sentences should be paraphrased to convey their intended meaning effectively. USER PROMPT: The English sentence is provided as follows: ¡english sentence¿ {sentence} ¡/english sentence¿ Please begin by assessing whether the English sentence contains any metaphors or similes. If there are none, respond with ”no metaphors and no similes.” If the English sentence does contain metaphors or similes, provide a literal translation of them, and then evaluate whether the literal translation is appropriate and easy for Chinese natives to understand. If it is suitable, format your response as follows (two lines): ”your literal translation for metaphors/similes here (in Chinese)” ”suitable” If it is unsuitable, please provide the reason for the unsuitability. Format your response as follows (three lines): ”your literal translation for metaphors/similes here (in Chinese)” ”unsuitable” ”reason for unsuitability here (in Chinese)”
@@ A.2 Prompts in Multi-Agent Framework
Given an English sentence, identify the important words (usually nouns, verbs, technical terms, and named entities that require special attention in translation) and translate them into Chinese. Output the translations in JSON format, for example: {”EnglishWord1”: ”ChineseTranslation”, ”EnglishWord2”: ”ChineseTranslation”} The Chinese translations can be a single translation or multiple options as deemed appropriate.
@@
SYSTEM PROMPT: Given an English sentence and a JSON object containing potential translations of important keywords, produce a Chinese literal translation of the entire sentence. Please directly output the Chinese translation without any descriptions. USER PROMPT: ¡English Sentence¿ {sentence} ¡/English Sentence¿ ¡Potential Keyword Translation¿ {keyword translation} ¡/Potential Keyword Translation¿
@@
SYSTEM PROMPT: Please evaluate the following Chinese translation of an English text. Rate the translation on a scale of 0 to 100, where: - 10 points: Poor translation; the text is somewhat understandable but contains significant errors and awkward phrasing that greatly hinder comprehension for a Chinese reader. - 30 points: Fair translation; the text conveys the basic meaning but lacks fluency and contains several awkward phrases or inaccuracies, making it challenging for a Chinese reader to fully grasp the intended message. - 50 points: Good translation; the text is mostly fluent and conveys the original meaning well, but may have minor awkwardness or slight inaccuracies that could confuse a Chinese reader. - 70 points: Very good translation; the text is smooth and natural, effectively conveying the intended meaning, but may still have minor issues that could slightly affect understanding for a Chinese reader. - 90 points: Excellent translation; the text is fluent and natural, conveying the original meaning clearly and effectively, with no significant issues that would hinder understanding for a Chinese reader. Please provide the reason first, followed by a score. Format your evaluation in the JSON structure below: {”reason”: ”reason for the score”, ”score”: int}
@@ A.3 Prompts in Thought Reformulation
A student is engaged in the task of translating an English sentence into Chinese. The English sentence is as follows: ¡English Sentence¿ {sentence} ¡/English Sentence¿ This student constantly thinks about and optimizes his translation. The whole process is shown as follows: ¡Translation Process¿ {translation process} ¡/Translation Process¿ Please polish the whole translation process into a long first-person self-reflection description (use the present tense). The self-reflection should begin with selecting the keywords from the English sentence, translating the keywords, and then attempt to translate the whole sentence, and then think about whether the translation is good or not, and iteratively make translation attempts. Finally, make a final translation decision. Output the self-reflection description directly without any additional descriptions or explanations. Each line in the self-reflection description can be regarded as a reasoning step to the translation.
@@ Appendix C GPT-4o Evaluator
Score the following translation from English to Chinese with respect to the human reference on a continuous scale from 0 to 100, where score of zero means ”no meaning preserved” and score of one hundred means ”perfect preservation of meaning, with faithfulness, expressiveness, and elegance”. English source: {src} Chinese human reference: {ref} Chinese translation: {hyp} Score:
@@
Score the following translation from English to Chinese on a continuous scale from 0 to 100, where score of zero means ”no meaning preserved” and score of one hundred means ”perfect preservation of meaning, with faithfulness, expressiveness, and elegance”. English source: {src} Chinese translation: {hyp} Score:
@@
DRT-8B $_{\text{(Backbone: Llama-3.1-8B-Instruct)}}$ 69.65 84.49 70.85 80.80 32.67 78.81
DRT-7B $_{\text{(Backbone: Qwen2.5-7B-Instruct)}}$ 75.05 85.57 71.78 82.38 35.54 80.19
DRT-14B $_{\text{(Backbone: Qwen2.5-14B-Instruct)}}$ 77.41 87.19 72.11 83.20 36.46 80.64
@@
[…]